# Distributional Reduction: Unifying Dimensionality Reduction and Clustering with Gromov-Wasserstein

### v2

@@ 3 Optimal Transport Formulation of DR
Gromov-Monge interpretation of DR. As suggested by eq. DR, dimension reduction seeks to find embeddings ${\mathbf{Z}}$ so that the similarity between the $(i,j)$ samples of the input data is as close as possible to the similarity between the $(i,j)$ samples of the embeddings. Under reasonable assumptions about ${\mathbf{C}}_{Z}$, this also amounts to identifying the embedding ${\mathbf{Z}}$ and the best permutation that realigns the two similarity matrices ${\mathbf{C}}_{X}({\mathbf{X}})$ and ${\mathbf{C}}_{Z}({\mathbf{Z}})$. Recall that the function ${\mathbf{C}}_{Z}$ is equivariant by permutation, if, for any $N\times N$ permutation matrix ${\mathbf{P}}$ and any ${\mathbf{Z}}$, ${\mathbf{C}}_{Z}({\mathbf{P}}{\mathbf{Z}})={\mathbf{P}}{\mathbf{C}}_{Z}({\mathbf{Z}}){\mathbf{P}}^{\top}$ bronstein2021geometric . This type of assumption is natural for ${\mathbf{C}}_{Z}$: if we rearrange the order of samples (i.e. the rows of ${\mathbf{Z}}$), we expect the similarity matrix between the samples to undergo the same rearrangement.
@@
Also, any solution ${\mathbf{Z}}$ of eq. DR is such that $({\mathbf{Z}},\operatorname{id})$ is solution of eq. 2. And conversely any $({\mathbf{Z}},\sigma)$ solution of eq. 2 is such that ${\mathbf{Z}}$ is a solution of eq. DR up to $\sigma$.
Lemma 3.1, proven in Section A.1, establishes an equivalence between eq. DR and the optimization of the embedding ${\mathbf{Z}}$ w.r.t a Gromov-Monge discrepancy [^2] Precisely $\min_{\sigma\in S_{N}}\sum_{ij}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[{\mathbf{C}}_{Z}({\mathbf{Z}})]_{\sigma(i)\sigma(j)})$ is the Gromov-Monge discrepancy between two discrete distributions, with the same number of atoms and uniform weights. memoli2018gromov given in eq. 2, which seeks to identify the permutation $\sigma$ that best aligns two similarity matrices, by solving a combinatorial quadratic assignment problem cela2013quadratic . We can delve deeper into these comparisons and demonstrate that the general formulation of dimension reduction is also equivalent to minimizing the Gromov-Wasserstein objective, which serves as a relaxation of the Gromov-Monge problem memoli2022comparison .
DR as GW Minimization. We suppose that the distributions have the same number of points ($N=n$) and uniform weights (${\mathbf{h}}_{Z}={\mathbf{h}}_{X}=\frac{1}{N}\bm{1}_{N}$). We recall that a matrix ${\mathbf{C}}\in\mathbb{R}^{N\times N}$ is conditionally positive definite (CPD), resp. conditionally negative definite (CND), if it is symmetric and $\forall{\mathbf{x}}\in\mathbb{R}^{N},{\mathbf{x}}^{\top}\mathbf{1}_{N}=0\text{s.t. }{\mathbf{x}}^{\top}{\mathbf{C}}{\mathbf{x}}\geq 0$, resp. $\leq 0$.
We thus extend Lemma 3.1 to the GW problem in the following (proof in Section A.2):
@@
The minimum of eq. DR is equal to $\min_{{\mathbf{Z}}}\operatorname{GW}_{L}({\mathbf{C}}_{X}({\mathbf{X}}),{\mathbf{C}}_{Z}({\mathbf{Z}}),\frac{1}{N}\bm{1}_{N},\frac{1}{N}\bm{1}_{N})$ if:
@@
    (spectral methods) ${\mathbf{C}}_{X}({\mathbf{X}})$ is any matrix, $L=L_{2}$ and ${\mathbf{C}}_{Z}({\mathbf{Z}})={\mathbf{Z}}{\mathbf{Z}}^{\top}$.
@@
    (neighbor embedding methods) $\operatorname{Im}({\mathbf{C}}_{X})\subseteq\mathbb{R}_{>0}^{N\times N}$, $L=L_{\operatorname{KL}}$, the matrix ${\mathbf{C}}_{X}({\mathbf{X}})$ is CPD and, for any ${\mathbf{Z}}$,
@@
    where $\bm{\alpha}_{\mathbf{Z}},\bm{\beta}_{\mathbf{Z}}\in\mathbb{R}^{N}_{>0}$ and ${\mathbf{K}}_{\mathbf{Z}}\in\mathbb{R}^{N\times N}_{>0}$ is such that $\log({\mathbf{K}}_{\mathbf{Z}})$ is CPD.
Remarkably, the first item of Theorem 3.2 shows that all spectral DR methods can be seen as OT problems in disguise, as they all equivalently minimize a GW problem. The second item also provides some insights into this equivalence in the case of neighbor embedding methods that require more assumptions. For instance, the Gaussian kernel ${\mathbf{K}}_{\mathbf{Z}}$, used extensively in DR (Section 2.1), satisfies the hypothesis as $\log({\mathbf{K}}_{\mathbf{Z}})=(-\|{\mathbf{z}}_{i}-{\mathbf{z}}_{j}\|_{2}^{2})_{ij}$ is CPD (see e.g. maron2018probably ). The terms $\bm{\alpha}_{\mathbf{Z}},\bm{\beta}_{\mathbf{Z}}$ also allow for considering all the usual normalizations of ${\mathbf{K}}_{\mathbf{Z}}$: by a scalar so as to have $\sum_{ij}[{\mathbf{C}}_{Z}({\mathbf{Z}})]_{ij}=1$, but also any row or doubly stochastic normalizations (with the Sinkhorn-Knopp algorithm sinkhorn1967concerning ).
Matrices satisfying $\log({\mathbf{K}}_{\mathbf{Z}})$ being CPD are well-studied in the literature and are known as infinitely divisible matrices bhatia2006infinitely . It is noteworthy that the t-Student kernel does not fall into this category. Moreover, in the aforementioned neighbor embedding methods, the matrix ${\mathbf{C}}_{X}({\mathbf{X}})$ is generally not CPD. The intriguing question of generalizing this result with weaker assumptions on ${\mathbf{C}}_{Z}$ and ${\mathbf{C}}_{X}$ remains open for future research. Interestingly, we have observed in the numerical experiments performed in Section 5 that the symmetric entropic affinity of van2023snekhorn was systematically CPD.
@@
In Corollary A.3 of the appendix we also provide other sufficient conditions for neighbor embedding methods with the cross-entropy loss $L(x,y)=x\log(x/y)$. They rely on specific structures for ${\mathbf{C}}_{Z}$ but do not impose any assumptions on ${\mathbf{C}}_{X}$. Additionally, in Section A.3, we provide a necessary condition based on a bilinear relaxation of the GW problem. Although its applicability is limited due to challenges in proving it in full generality, it requires minimal assumptions on ${\mathbf{C}}_{X},{\mathbf{C}}_{Z}$ and $L$.
@@ 4.1 Distributional Reduction Problem
$\min_{{\mathbf{Z}}\in\mathbb{R}^{n\times d}}\min_{{\mathbf{h}}_{Z}\in\Sigma_{n}}\operatorname{GW}_{L}({\mathbf{C}}_{X}({\mathbf{X}}),{\mathbf{C}}_{Z}({\mathbf{Z}}),{\mathbf{h}}_{X},{\mathbf{h}}_{Z})$ (DistR)
@@
Figure 2: GW projections of a single-cell genomics dataset chen2019high on regular grids with increasing resolutions, respectively encoded as $C_{X}({\mathbf{X}})={\mathbf{X}}{\mathbf{X}}^{\top}$ and $C_{Z}({\mathbf{Z}})={\mathbf{Z}}{\mathbf{Z}}^{\top}$. Pixels on cropped grids are colored by interpolating samples’ colors according to the transport plan and their intensity is proportional to their mass.
@@
For a given embedding ${\mathbf{Z}}$ and $L=L_{2}$, it is known that minimizing the DistR objective w.r.t ${\mathbf{h}}_{Z}$ is equivalent to a problem that is computationally simpler than the usual GW one, namely the semi-relaxed GW divergence $\operatorname{srGW}_{L}$ vincent2021semi :
@@
where ${\mathcal{U}}_{n}({\mathbf{h}}_{X}):=\left\{{\mathbf{T}}\in\mathbb{R}_{+}^{N\times n}:{\mathbf{T}}\bm{1}_{n}={\mathbf{h}}_{X}\right\}$. To efficiently address eq. DistR, we first observe that this equivalence holds for any inner divergence $L$ with a straightforward adaptation of the proof in vincent2021semi . Additionally, we prove that $\operatorname{srGW}_{L}$ remains a divergence as soon as $L$ is itself a divergence. Consequently, $\operatorname{srGW}_{L}$ vanishes iff both compared measures are isomorphic in a weak sense chowdhury2019gromov . We emphasize that taking a proper divergence $L$ is important (and basic assumptions on ${\mathbf{X}}$), as it avoids some trivial solutions as detailed in Appendix B.
@@
Computation. DistR is a non-convex problem that we propose to tackle using a Block Coordinate Descent algorithm (BCD, tseng2001convergence ) guaranteed to converge to local optimum grippo2000convergence ; Lyu2023bmm . The BCD alternates between the two following steps. First, we optimize in ${\mathbf{Z}}$ for a fixed transport plan using gradient descent with adaptive learning rates kingma2014adam . Then we solve for a srGW problem given ${\mathbf{Z}}$. To this end, we extended both solvers proposed in vincent2021semi to support losses $L_{2}$ and $L_{\operatorname{KL}}$. Namely, the Mirror Descent promoting smoothness along iterations depending on a hyparameter $\varepsilon>0$ and the Conditional Gradient seen as the edge-case $\varepsilon=0$. Following Proposition 1 in peyre2016gromov , a vanilla implementation leads to $\mathcal{O}(nN^{2}+n^{2}N)$ operations to compute the loss or its gradient. However in many DR methods, ${\mathbf{C}}_{X}({\mathbf{X}})$ or ${\mathbf{C}}_{Z}({\mathbf{Z}})$, or their transformations within the loss $L$, admit explicit low-rank factorizations. Including e.g. matrices involved in spectral methods and other similarity matrices derived from squared Euclidean distance matrices scetbon2022linear . In these settings, we exploit these factorizations to reduce the computational complexity of our solvers down to $\mathcal{O}(Nn(p+d)+(N+n)pd+min(n^{2},nd^{2}))$ when $L=L_{2}$, and $O(Nnd+n^{2}d)$ when $L=L_{\operatorname{KL}}$. We refer the reader interested in these algorithmic details to Appendix D. As shown in Section E.6, DistR runs in a few seconds on datasets with several thousands of samples.
Related work. The CO-Optimal Transport (COOT) clustering approach proposed in redko2020co is the closest to our work. As detailed in Section E.2, it simultaneously estimates sample and feature clustering and thus can be directly applied to perform joint clustering and DR. Despite SOTA performances on co-clustering tasks, COOT-clustering corresponds to a linear DR method which is limiting for joint clustering and DR tasks. In Section E.2, we show that COOT leads to a low average homogeneity of the prototypes $\{{\mathbf{z}}_{k}\}_{k\in{[\![n]\!]}}$, indicating that it often assigns points with different labels to the same prototype. In contrast, DistR exploits the more expressive non-linear similarity functions offered by existing DR methods and leads to significantly improved results. Other joint DR-clustering approaches, such as liu2022joint , involve modeling latent variables by a mixture of distributions. In comparison, DistR is more versatile, as it easily adapts to any $(L,{\mathbf{C}}_{X},{\mathbf{C}}_{Z})$ of existing DR methods.
@@ 4.2 Clustering Properties
We elaborate now on the links between DistR and clustering methods. In what follows, we call a coupling ${\mathbf{T}}\in[0,1]^{N\times n}$ with a single non-null element per row a membership matrix. When the coupling is a membership matrix each data point is associated with a single prototype thus achieving a hard clustering of the input samples. We will see that a link can be drawn with kernel k-means using the analogy of GW barycenters. More precisely the srGW barycenter vincent2021semi seeks for a closest target graph $(\overline{{\mathbf{C}}},\overline{{\mathbf{h}}})$ from $({\mathbf{C}}_{X},{\mathbf{h}}_{X})$ by solving
@@
We establish below that srGWB is of particular interest for clustering. The motivation for this arises from the findings of chen2023gromov , which demonstrate that, when ${\mathbf{C}}_{X}({\mathbf{X}})$ is positive semi-definite and ${\mathbf{T}}$ is constrained to belong to the set of membership matrices (as opposed to couplings in ${\mathcal{U}}_{n}({\mathbf{h}})$), eq. srGWB is equivalent to a kernel k-means whose samples are weighted by ${\mathbf{h}}_{X}$ dhillon2004kernel ; dhillon2007weighted . Interestingly, these additional constraints are unnecessary. Indeed, we show below that the original srGWB problem admits membership matrices as the optimal coupling for a broader class of ${\mathbf{C}}_{X}({\mathbf{X}})$ input matrices (see proof in Appendix C).
@@
Let ${\mathbf{h}}_{X}\in\Sigma_{N}$ and $L=L_{2}$. Suppose that for any ${\mathbf{X}}\in\mathbb{R}^{N\times p}$ the matrix ${\mathbf{C}}_{X}({\mathbf{X}})$ is CPD or CND. Then the problem eq. srGWB admits a membership matrix as optimal coupling.
Theorem 4.1 and relations proven in chen2023gromov state that eq. srGWB is equivalent to the aforementioned kernel k-means when ${\mathbf{C}}_{X}({\mathbf{X}})$ is positive semi-definite. Moreover, as the (hard) clustering property holds for more generic types of matrices, namely CPD and CND, srGWB stands out as a fully-fledged clustering method. Although these results do not apply directly to DistR, except if e.g. the dimension $d$ is set to the unknown rank of the srGW barycenter, we argue that they further legitimize the use of GW projections for clustering. Interestingly, we also observe in practice that the couplings obtained by DistR are always membership matrices, regardless of ${\mathbf{C}}_{Z}$. Further research need to be carried out to better understand this observation.

### v3

@@ 3 OT Formulation of DR
Gromov-Monge interpretation of DR. As suggested by eq. DR, dimension reduction seeks to find embeddings ${\mathbf{Z}}$ so that the similarity between the $(i,j)$ samples of the input data is as close as possible to the similarity between the $(i,j)$ samples of the embeddings. Under equivariant assumptions on ${\mathbf{C}}_{Z}$, this also amounts to identifying the embedding ${\mathbf{Z}}$ and the best permutation that realigns the two similarity matrices ${\mathbf{C}}_{X}({\mathbf{X}})$ and ${\mathbf{C}}_{Z}({\mathbf{Z}})$. Recall that the function ${\mathbf{C}}_{Z}$ is equivariant by permutation, if, for any $N\times N$ permutation matrix ${\mathbf{P}}$ and any ${\mathbf{Z}}$, ${\mathbf{C}}_{Z}({\mathbf{P}}{\mathbf{Z}})={\mathbf{P}}{\mathbf{C}}_{Z}({\mathbf{Z}}){\mathbf{P}}^{\top}$ (Bronstein et al., 2021). This type of assumption is natural for ${\mathbf{C}}_{Z}$: if we rearrange the order of samples (i.e. the rows of ${\mathbf{Z}}$), we expect the similarity matrix between the samples to undergo the same rearrangement.
@@
Also, any sol. ${\mathbf{Z}}$ of eq. DR is such that $({\mathbf{Z}},\operatorname{id})$ is solution of eq. 2. And conversely any $({\mathbf{Z}},\sigma)$ sol. of eq. 2 is such that ${\mathbf{Z}}$ is a solution of eq. DR up to $\sigma$.
Lemma 3.1, proven in Section B.1, establishes an equivalence between eq. DR and the optimization of the embedding ${\mathbf{Z}}$ w.r.t a Gromov-Monge discrepancy (Mémoli and Needham, 2018) given in eq. 2, which seeks to identify the permutation $\sigma$ that best aligns two similarity matrices, by solving a combinatorial quadratic assignment problem (Cela, 2013). We can delve deeper into these comparisons and demonstrate that the general formulation of dimension reduction is also equivalent to minimizing the Gromov-Wasserstein objective, which serves as a relaxation of the Gromov-Monge problem (Mémoli and Needham, 2022).
DR as GW Minimization. We suppose that the distributions have the same number of points ($N=n$) and uniform weights (${\mathbf{h}}_{Z}={\mathbf{h}}_{X}=\frac{1}{N}\bm{1}_{N}$). We recall that a matrix ${\mathbf{C}}\in\mathbb{R}^{N\times N}$ is conditionally positive definite (CPD), resp. conditionally negative definite (CND), if it is symmetric and $\forall{\mathbf{x}}\in\mathbb{R}^{N}\text{ s.t. }{\mathbf{x}}^{\top}\mathbf{1}_{N}=0\text{ we have }{\mathbf{x}}^{\top}{\mathbf{C}}{\mathbf{x}}\geq 0$, resp. $\leq 0$.
We thus extend Lemma 3.1 to the GW problem in the following (proof in Section B.2):
@@
The minimum eq. DR is equal to $\min_{{\mathbf{Z}}}\operatorname{GW}_{L}({\mathbf{C}}_{X}({\mathbf{X}}),{\mathbf{C}}_{Z}({\mathbf{Z}}),\frac{1}{N}\bm{1}_{N},\frac{1}{N}\bm{1}_{N})$ in the following cases:
@@
    (spectral methods) When ${\mathbf{C}}_{X}({\mathbf{X}})$ is any matrix, $L=L_{2}$ and ${\mathbf{C}}_{Z}({\mathbf{Z}})={\mathbf{Z}}{\mathbf{Z}}^{\top}$.
@@
    (neighbor embedding methods) When $\operatorname{Im}({\mathbf{C}}_{X})\subseteq\mathbb{R}_{>0}^{N\times N}$, $L=L_{\operatorname{KL}}$, the matrix ${\mathbf{C}}_{X}({\mathbf{X}})$ is CPD and, for any ${\mathbf{Z}}$,
@@
    for some $\bm{\alpha}_{\mathbf{Z}},\bm{\beta}_{\mathbf{Z}}\in\mathbb{R}^{N}_{>0}$ and ${\mathbf{K}}_{\mathbf{Z}}\in\mathbb{R}^{N\times N}_{>0}$ such that $\log({\mathbf{K}}_{\mathbf{Z}})$ is CPD.
Remarkably, the first item of Theorem 3.2 shows that all spectral DR methods can be seen as OT problems in disguise, as they all equivalently minimize a GW problem. The second item also provides some insights into this equivalence in the case of neighbor embedding methods that require more assumptions. For instance, the Gaussian kernel ${\mathbf{K}}_{\mathbf{Z}}$, used extensively in DR (Section 2.1), satisfies the hypothesis as $\log({\mathbf{K}}_{\mathbf{Z}})=(-\|{\mathbf{z}}_{i}-{\mathbf{z}}_{j}\|_{2}^{2})_{ij}$ is CPD (see e.g. Maron and Lipman 2018). The terms $\bm{\alpha}_{\mathbf{Z}},\bm{\beta}_{\mathbf{Z}}$ also allow for considering all the usual normalizations of ${\mathbf{K}}_{\mathbf{Z}}$: by a scalar so as to have $\sum_{ij}[{\mathbf{C}}_{Z}({\mathbf{Z}})]_{ij}=1$, but also any row or doubly stochastic normalizations (with the Sinkhorn-Knopp algorithm Sinkhorn and Knopp 1967).
Matrices satisfying $\log({\mathbf{K}}_{\mathbf{Z}})$ being CPD are well-studied in the literature and are known as infinitely divisible matrices (Bhatia, 2006). It is noteworthy that the t-Student kernel does not fall into this category. Moreover, in the aforementioned neighbor embedding methods, the matrix ${\mathbf{C}}_{X}({\mathbf{X}})$ is generally not CPD. The intriguing question of generalizing this result with weaker assumptions on ${\mathbf{C}}_{Z}$ and ${\mathbf{C}}_{X}$ remains open for future research. Interestingly, we have observed in the numerical experiments performed in Section 5 that the symmetric entropic affinity of Van Assel et al. (2023) was systematically CPD.
@@
In Corollary B.3 of the appendix we also provide other sufficient conditions for neighbor embedding methods with the cross-entropy loss $L(x,y)=x\log(x/y)$. They rely on specific structures for ${\mathbf{C}}_{Z}$ but do not impose any assumptions on ${\mathbf{C}}_{X}$. Additionally, in Section B.3, we provide a necessary condition based on a bilinear relaxation of the GW problem. Although its applicability is limited due to challenges in proving it in full generality, it requires minimal assumptions on ${\mathbf{C}}_{X},{\mathbf{C}}_{Z}$ and $L$.
@@ 4.1 Distributional Reduction Problem
$\min_{\begin{smallmatrix}{\mathbf{Z}}\in\mathbb{R}^{n\times d}\\ {\mathbf{h}}_{Z}\in\Sigma_{n}\end{smallmatrix}}\operatorname{GW}_{L}({\mathbf{C}}_{X}({\mathbf{X}}),{\mathbf{C}}_{Z}({\mathbf{Z}}),{\mathbf{h}}_{X},{\mathbf{h}}_{Z})\,.$ (DistR)
@@
Figure 2: GW projections of a genomics dataset (Chen et al., 2019) on regular grids with increasing resolutions, respectively encoded as $C_{X}({\mathbf{X}})={\mathbf{X}}{\mathbf{X}}^{\top}$ and $C_{Z}({\mathbf{Z}})={\mathbf{Z}}{\mathbf{Z}}^{\top}$. Pixels on cropped grids are colored by interpolating samples’ colors according to the transport plan and their intensity is proportional to their mass.
@@
For a given embedding ${\mathbf{Z}}$ and $L=L_{2}$, it is known that minimizing the DistR objective w.r.t ${\mathbf{h}}_{Z}$ is equivalent to a problem that is computationally simpler than the usual GW one, namely the semi-relaxed GW divergence $\operatorname{srGW}_{L}$ (Vincent-Cuaz et al., 2022a) given by
@@
where ${\mathcal{U}}_{n}({\mathbf{h}}_{X}):=\left\{{\mathbf{T}}\in\mathbb{R}_{+}^{N\times n}:{\mathbf{T}}\bm{1}_{n}={\mathbf{h}}_{X}\right\}$. To efficiently address eq. DistR, we first observe that this equivalence holds for any inner divergence $L$. Additionally, we prove that $\operatorname{srGW}_{L}$ remains a divergence as soon as $L$ is itself a divergence. Consequently, $\operatorname{srGW}_{L}$ vanishes iff both compared measures are isomorphic in a weak sense (Chowdhury and Mémoli, 2019). We emphasize that taking a proper divergence $L$ is important (and basic assumptions on ${\mathbf{X}}$), as it avoids some trivial solutions. These results are detailed in Appendix C.
@@
Computation. DistR is a non-convex problem that we propose to tackle using a Block Coordinate Descent algorithm (BCD, Tseng 2001) guaranteed to converge to local optimum (Grippo and Sciandrone, 2000, Lyu and Li, 2023). The BCD alternates between the two following steps. First, we optimize in ${\mathbf{Z}}$ for a fixed transport plan using gradient descent with adaptive learning rates (Kingma and Ba, 2014). Then we solve for a srGW problem given ${\mathbf{Z}}$. To this end, we extended both solvers proposed in Vincent-Cuaz et al. (2022a) to support losses $L_{2}$, $L_{\operatorname{KL}}$ and $L_{BCE}$ (see Section E.1). Namely, the Mirror Descent promoting smoothness along iterations depending on a hyparameter $\varepsilon>0$ and the Conditional Gradient seen as the edge-case $\varepsilon=0$. Following Proposition 1 in (Peyré et al., 2016), a vanilla implementation leads to $\mathcal{O}(nN^{2}+n^{2}N)$ operations to compute the loss or its gradient. However in many DR methods, ${\mathbf{C}}_{X}({\mathbf{X}})$ or ${\mathbf{C}}_{Z}({\mathbf{Z}})$, or their transformations within the loss $L$, admit explicit low-rank factorizations. Including e.g. matrices involved in spectral methods and other similarity matrices derived from squared Euclidean distance matrices (Scetbon et al., 2022). In these settings, we exploit these factorizations to reduce the computational complexity of our solvers down to $\mathcal{O}(Nn(p+d)+(N+n)pd+min(n^{2},nd^{2}))$ when $L=L_{2}$, and $O(Nnd+n^{2}d)$ when $L=L_{\operatorname{KL}}$. We refer the reader interested in these algorithmic details to Appendix E. As shown in Section F.9, DistR already runs in a few seconds on datasets with several thousands of samples without leveraging low-rank properties.
Related work. The CO-Optimal Transport (COOT) clustering approach proposed in Redko et al. (2020) is the closest to our work. It simultaneously estimates sample and feature clustering and thus can be directly applied to perform joint clustering and DR. Despite SOTA performances on co-clustering tasks, COOT-clustering corresponds to a linear DR method which is limiting for joint clustering and DR tasks. In Section F.2, we show that COOT leads to a low average homogeneity of the prototypes $\{{\mathbf{z}}_{k}\}_{k\in{[\![n]\!]}}$, indicating that it often assigns points with different labels to the same prototype. In contrast, DistR exploits the more expressive non-linear similarity functions offered by existing DR methods and leads to significantly improved results. Other joint DR-clustering approaches, such as Liu et al. (2022a), involve modeling latent variables by a mixture of distributions. In comparison, DistR is more versatile, as it easily adapts to any $(L,{\mathbf{C}}_{X},{\mathbf{C}}_{Z})$ of existing DR methods.
@@ 4.2 Clustering Properties
We elaborate now on the links between DistR and clustering methods. In what follows, we call a coupling ${\mathbf{T}}\in[0,1]^{N\times n}$ with a single non-null element per row a membership matrix. When the coupling is a membership matrix each data point is associated with a single prototype thus achieving a hard clustering of the input samples. We will see that a link can be drawn with kernel k-means using the analogy of GW barycenters. More precisely the srGW barycenter (Vincent-Cuaz et al., 2022a) seeks for a closest target graph $(\overline{{\mathbf{C}}},\overline{{\mathbf{h}}})$ from $({\mathbf{C}}_{X},{\mathbf{h}}_{X})$ by solving
@@
We establish below that srGWB is of particular interest for clustering. The motivation for this arises from the findings of (Chen et al., 2023), which demonstrate that, when ${\mathbf{C}}_{X}({\mathbf{X}})$ is positive semi-definite and ${\mathbf{T}}$ is constrained to belong to the set of membership matrices (as opposed to couplings in ${\mathcal{U}}_{n}({\mathbf{h}})$), eq. srGWB is equivalent to a kernel k-means whose samples are weighted by ${\mathbf{h}}_{X}$ (Dhillon et al., 2004, 2007). Interestingly, these additional constraints are unnecessary. Indeed, we show below that the original srGWB problem admits membership matrices as the optimal coupling for a broader class of ${\mathbf{C}}_{X}({\mathbf{X}})$ input matrices for $L=L_{2}$ (see proof in Appendix D).
@@
Let ${\mathbf{h}}_{X}\in\Sigma_{N}$ and $L=L_{2}$. Suppose that for any ${\mathbf{X}}\in\mathbb{R}^{N\times p}$ the matrix ${\mathbf{C}}_{X}({\mathbf{X}})$ is CPD or CND. Then the problem eq. srGWB admits a membership matrix as optimal coupling, i.e. , there is a minimizer of ${\mathbf{T}}\in\mathcal{U}_{n}({\mathbf{h}}_{X})\to\min_{\overline{{\mathbf{C}}}\in\mathbb{R}^{n\times n}}E_{L}({\mathbf{C}}_{X}({\mathbf{X}}),\overline{{\mathbf{C}}},{\mathbf{T}})$ with only one non-zero value per row.
The implications of this theorem are as follows. First, as shown in Appendix D, given an optimal plan ${\mathbf{T}}$, the coefficient $(i,j)$ of the barycenter $\overline{{\mathbf{C}}}$ can be written as $\frac{1}{n_{i}n_{j}}\sum_{pq}[{\mathbf{C}}_{X}({\mathbf{X}})]_{pq}T_{pi}T_{qj}$ where $n_{i}=\sum_{p}T_{pi}$. Consequently, when ${\mathbf{T}}$ is a membership matrix, $\overline{{\mathbf{C}}}$ represent graph weights, where each element $(i,j)$ is the weighted sum of the original graph weights ${\mathbf{C}}_{X}({\mathbf{X}})$ for nodes in the clusters of $i$ and $j$, following standard edge contraction methods (Loukas, 2019). Second, Theorem 4.1 and relations proven in Chen et al. (2023) state that eq. srGWB is equivalent to the aforementioned kernel k-means when ${\mathbf{C}}_{X}({\mathbf{X}})$ is positive semi-definite. This equivalence also emphasizes spectrum-preservating properties of the srGWB problem, whose optimal value satisfies $E_{L_{2}}({\mathbf{C}}_{X}({\mathbf{X}}),\overline{{\mathbf{C}}},{\mathbf{T}})=\sum_{i=1}^{N}\lambda_{i}^{2}-\sum_{j=1}^{n}\nu_{j}^{2}$, where $\{\lambda_{i}\}_{i=1}^{N}$ and $\{\nu_{j}\}_{j=1}^{n}$ are the eigenvalues sorted in descending order of $\text{diag}({\mathbf{h}}_{X})^{-1/2}{\mathbf{C}}_{X}({\mathbf{X}})\text{diag}({\mathbf{h}}_{X})^{-1/2}$ and $\overline{{\mathbf{C}}}$ respectively, and satisfy the Pointcaré Separation Theorem. Finally, as the (hard) clustering property holds for more generic types of matrices, namely CPD and CND, srGWB stands out as a fully-fledged clustering method. Although these results do not apply directly to DistR, except if e.g. the dimension $d$ is set to the unknown rank of the srGW barycenter, we argue that they further legitimize the use of GW projections for clustering. Interestingly, we also observe in practice that the couplings obtained by DistR are always membership matrices, regardless of ${\mathbf{C}}_{Z}$.